whitemonsterbunny17 (whitemonsterbunny17):

Can someone please help me with simplifying square roots? Pretty please, with a rainbow cupcake on top? :3 Any assistance is greatly appreciated! <3

whitemonsterbunny17 (whitemonsterbunny17):

\(\Large\bf\color{blueviolet}{Simplify:}\) \(\Large\sqrt{72}\)

whitemonsterbunny17 (whitemonsterbunny17):

I really don't get the concept of square roots in general, but I know that to simplify, you need to take out something called "perfect squares"?

OpenStudy (phi):

factor the 72 into primes, and look for pairs of numbers that you can "take out"

OpenStudy (phi):

a simple example \[ \sqrt{4} = \sqrt{2\cdot 2} = 2 \] or \[ \sqrt{8} = \sqrt{2\cdot 2\cdot 2} = 2\sqrt{2} \] this last example has a pair of 2's that come out, and 1 2 left over that stays inside.

OpenStudy (phi):

for 72, I would divide by 2 to get 2*36 now divide the 36 by 2 to get 2*2*18 keep going ?

whitemonsterbunny17 (whitemonsterbunny17):

I can somewhat get it. I get how you got 36 and 18.

OpenStudy (phi):

I would learn (some) of these divisibility rules https://en.wikipedia.org/wiki/Divisibility_rule#Divisibility_rules_for_numbers_1.E2.80.9320 I would learn as many of the primes as you can: http://metricconversion.biz/list-of-first-100-prime-numbers.html

whitemonsterbunny17 (whitemonsterbunny17):

So, each time you divide, you just add a 2 to multiply by, correct?

OpenStudy (phi):

do you agree that 2*36 = 72 ? now look at 36. it is even, so it is divisible by 2: and 36= 2*18 that means 2*36 can be written as 2*(2*18) or 2*2*18 (and if you multiply it out you get 72) now 18 is even, so we can divide it by 2: 18 = 2*9 replace 18 with 2*9 and our list is now 2*2*2*9

OpenStudy (phi):

9 is divisible by 3 : 9= 3*3 so the list is 2*2*2*3*3 if you multiply that out , you get 72 2 and 3 are prime (nothing divides into them so we are done factoring now circle pairs of the same number.

whitemonsterbunny17 (whitemonsterbunny17):

Okay, I'm starting to get it.. So, would it be 2^3*3^2 ?

OpenStudy (phi):

yes you can write it like that (easier, especially for 2^10 , for example) but for these problems, I would stick with 2*2*2*3*3 and circle one pair of 2's and one pair of 3's (leaving a unmatched 2 , which stays inside the square root)

whitemonsterbunny17 (whitemonsterbunny17):

Okay, but there's no square root symbol at the end, though. o-O This is how my math teacher showed it (which I wasn't getting) \(\sqrt{12}\) = \(\sqrt{4}\) \(\sqrt{3}\) = \(\large2\) \(\sqrt{3}\)

OpenStudy (phi):